A stretched wire has a Young’s modulus Y and energy density E. the strain in the stretched wire is ?

Grade:11

2 Answers

Arun
3 years ago

As we discussed in concept

Stress -

The internal restoring force per unit area of deformed body.

- wherein

Due to the forces, the body gets deformed and internal forces appear.

 

 

Stress -

Stress =\frac{force}{Area}=\frac{F}{A}
 

- wherein

F= applied \: force

A= area \:

 

 E=\frac{1}{2}\times Stress \times Strain

E=\frac{1}{2}\times Stress \times \frac{Stress}{Y}\:=\:\frac{1}{2}\:\frac{S^{2}}{Y}

E=\frac{1}{2Y}\:Y^{2}(Strain)^{2}\:\:\:\:\:\:\:\:\:\:\:\Rightarrow 2E=YS^{2}

S=\sqrt{\frac{2E}{Y}}


 
Vikas TU
3 years ago
Dear stuent 
A stretched wire has a Young’s modulus Y and energy density E. the strain in the stretched wire is  sqrt(2E/Y).